OpenStudy (anonymous):

Find the area of a 120° sector of a circle whose radius is 6.

jimthompson5910 (jim_thompson5910):

Area of a full circle A = pi*r^2

jimthompson5910 (jim_thompson5910):

Area of a sector of a circle A = (angle/360)*(Area of full circle) A = (angle/360)*pi*r^2

OpenStudy (anonymous):

answer has to be in pi

OpenStudy (anonymous):

339.12

jimthompson5910 (jim_thompson5910):

A = (angle/360)*pi*r^2 A = (120/360)*pi*6^2 Remember to leave the 'pi' alone (and don't replace it with 3.14...)

OpenStudy (anonymous):

so it would be 3 x 36 = 108

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

120/360 isn't 3 though

jimthompson5910 (jim_thompson5910):

you divided in the wrong order

OpenStudy (anonymous):

.33

OpenStudy (anonymous):

.33 x 36 11.88

jimthompson5910 (jim_thompson5910):

it's more like 0.333333 where the 3's repeat forever

jimthompson5910 (jim_thompson5910):

or just 1/3

jimthompson5910 (jim_thompson5910):

(1/3)*(36) = 36/3 = 12 So the answer is 12pi

OpenStudy (anonymous):

so the answer is 12

OpenStudy (anonymous):

thanks for helping

jimthompson5910 (jim_thompson5910):

12pi, but yes you pretty much have it

jimthompson5910 (jim_thompson5910):

yw